AnxiTEA:

help please? C= 5 9 (F−32) The equation above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of 5 9 degree Celsius. A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. A temperature increase of 5 9 degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius. A) I only B) II only C) III only D) I and II only

Angle:

Fractions don't work out well when copy pasting them do you have a picture that you can attach?

AnxiTEA:

im on computer-

Angle:

Let's play with a few numbers \(C=\frac{5}{9}(F-32)\) If we plug in F = 41 into the formula: \(C=\frac{5}{9}(F-32)=\frac{5}{9}(41-32)=\frac{5}{9}(9)=5\) Let's test out your options: If we increase F by 1, the new F = 42 \(C=\frac{5}{9}(F-32)=\frac{5}{9}(42-32)=\frac{5}{9}(10)=5.555...\) The temperature increased by 0.55555 = 5/9 when we increased F by 1 does that make your option I true?

AnxiTEA:

i found the answer